OpenStudy (anonymous):

How many different 12-member juries can be chosen from a pool of 45 people?

OpenStudy (anonymous):

you are asked to compute \[\binom{45}{12}\] or \[_{45}C_{12}\]

OpenStudy (anonymous):

i would use a calculator, is a rather large and annoying number to compute

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=45+choose+12

OpenStudy (anonymous):

if you have a lot of patients you can compute \[\frac{45\times44\times43\times42\times41\times40\times39\times38\times37\times36\times35\times34 }{12\times 11\times 10\times 9\times 8\times 7\times 6\times 5\times 4\times 3\times 2}\]

OpenStudy (anonymous):

or even *patience

OpenStudy (anonymous):

Thank you satellite73!